# Schema-based active inference supports rapid generalization of experience and frontal cortical coding of abstract structure

### v1

@@ Abstract
Schemas – abstract relational structures that capture the commonalities across experiences – are thought to underlie humans’ and animals’ ability to rapidly generalize knowledge, rebind new experiences to existing structures, and flexibly adapt behavior across contexts. Despite their central role in cognition, the computational principles and neural mechanisms supporting schema formation and use remain elusive. Here, we introduce schema-based hierarchical active inference (S-HAI), a novel computational framework that combines predictive processing and active inference with schema-based mechanisms. In S-HAI, a higher-level generative model encodes abstract task structure, while a lower-level model encodes spatial navigation, with the two levels linked by a grounding likelihood that maps abstract goals to physical locations. Through a series of simulations, we show that S-HAI reproduces key behavioral signatures of rapid schema-based generalization in spatial navigation tasks, including the ability to flexibly remap abstract schemas onto novel contexts, resolve goal ambiguity, and balance reuse versus accommodation of novel mappings. Crucially, S-HAI also reproduces prominent neural codes reported in rodent medial prefrontal cortex during a schema-dependent navigation and decision task, including task-invariant goal-progress cells, goal-identity cells, and goal-and-spatially conjunctive cells, as well as place-like codes at the lower level. Taken together, these results provide a mechanistic account of schema-based learning and inference that bridges behavior, neural data, and theory. More broadly, our findings suggest that schema formation and generalization may arise from predictive processing principles implemented hierarchically across cortical and hippocampal circuits, enabling the generalization of experience.
@@ 1 Introduction
Finally, various studies indicate that the frontal cortex—possibly in interaction with the hippocampus—plays a crucial role in forming cognitive maps and schematic associations, integrating new information into existing schemas, and determining whether assimilation or accommodation processes are engaged (Gilboa and Marlatte, 2017; Zeithamova et al., 2012; Van Kesteren et al., 2012; Giuliano et al., 2021; Bonasia et al., 2018; Baldassano et al., 2018; Basu et al., 2021; Manakov et al., 2025; Schuck et al., 2016; Wang and Hayden, 2021; Vaidya and Badre, 2022; Tang et al., 2023). A particularly striking example of schema learning and generalization is provided by El-Gaby et al. (2024), who demonstrated that the rodent medial frontal cortex is involved in learning the abstract structure of a sequential task—the ABCD task, which requires reaching four goal locations in the correct order—and in reusing this structure in novel environments where the sequence remains the same but the goal locations vary. A key indicator of fast, schema-based learning is that rodents rapidly move toward location A after discovering location D. At the neuronal level, this ability is supported by various types of cells that are sensitive to different combinations of abstract goal, physical location, and other task-relevant information.
@@
While prior work on CSCGs has addressed the problem of learning space as a latent structure from partially observable sensations (Raju et al., 2024), and CSCG schemas (Guntupalli et al., 2023; Swaminathan et al., 2023) addresses the problem of transferring a learned spatial or algorithm structure using schemas, they do not address the problem of transferring multiple task structures that are learned in the same spatial environment. To do this in a partially observable setting, new task-structure schemas have to learned that uses the latent states of the spatial structure that was previously learned.
@@ 2.2 Schema-based hierarchical active inference (S-HAI)
Crucially, the S-HAI agent also includes a grounding likelihood: a probabilistic mapping between the abstract schema representing transitions between goals in task space (i.e., A, B, C, and D) and the concrete locations of the goals in navigation space (i.e., the locations in the maze where it can find reward). Learning grounding likelihoods is what affords schema-based generalization: it enables the agent to rapidly map its abstract schema (e.g., ABCD) to each novel spatial configuration of the goals (i.e., each block), rather than relearning the correct sequence of actions and goals from scratch on each trial. In our simulations, we implement online learning of the grounding likelihood, alongside both an online and offline variant for the Level 2 schema. In addition, we introduce a mixture model over grounding likelihoods that allows the S-HAI agent to flexibly infer which of its existing grounding likelihoods is most useful in the current maze, or to create a new one when needed. Previewing our results, we found that in all cases (online or offline, with or without mixture), grounding likelihoods allow the S-HAI agent to outperform alternative models trained on the same or even larger datasets but without schema learning. Learning the grounding likelihoods is similar to learning the emission matrix of a schema as in (Guntupalli et al., 2023) and (Swaminathan et al., 2023).
@@
In the hierarchical active inference (HAI) agent without schemas, Level 2 is implemented using a clone-structured graph (CSCG) (George et al., 2021) and the grounding likelihood is an identity matrix. The agent is denoted as HAI-$i$, with the index $i$ indicating the number of tasks the agent is trained on; training is done offline on a sequence of the first $i$ tasks, with 10 000 interaction steps per task. See Figure 1 for the transition dynamics between goals learned by the HAI-$2$ agent (trained on two blocks) at Level 2. Note that, unlike the S-HAI agent, which learns a unique sequential schema, the HAI-$2$ agent learns a unique subcycle for each block. When the HAI-$i$ agent is trained on more blocks, it tends to learn multiple block-specific subcycles (not shown here, but similar to what is reported in (Van de Maele et al., 2024)).
@@
Figure 2 shows the simulation results for the ABCD task, reporting the average reward rate (smoothed over 250 steps), normalized against optimal performance, across 40 blocks. The solid line denotes the mean across blocks, and shaded areas indicate the standard error. Each block ended when the agent reached 10 000 interactions with the environment. As expected, the HAI agents without schemas (trained offline on half (20) or all (40) blocks; HAI-20, and HAI-40) outperformed the Random baseline, with performance improving as the number of training blocks increased. However, the HAI-20 agent trained in half of the blocks showed limited generalization to novel blocks beyond its training set.
@@
Figure 4: Simulation of neural activity in the S-HAI model during the ABCD task. (a) Four blocks of the ABCD task. (b) Simulated activity of “goal-progress” cells during the ABCD task, plotted in a circle. The nodes represent locations visited during the task and their colors nodes indicate the normalized expected inductive cost, under the variational posterior over the level 1 state; these range from purple (low expectation, corresponding to early phase of progress towards a goal) to green (mid expectation and mid phase of progress) and yellow (high expectation and late phase of progress). The line indicates where the reward was received, and the color indicates which phase of the task this was in. (c) Simulated activity of “goal-progress” cells during the ABCD task, plotted in space. The panels show activations of the normalized expected inductive cost under the variational posterior over level 1 state in distinct phases of Task 4. As before, purple, green, and yellow represent low expectation (early phase), mid expectation (mid phase), and high expectation (late phase of progress). Note that in the third panel, there are two paths from C to D, i.e., the animal can move upward or to the left after leaving C. (d) Level 2 state activations during the last 1000 steps in each environment for the tasks. Dots are plotted at points between two visited tiles, colored by the high-level state. Locations are augmented with random noise for visual clarity. (e) Dots colored according to the joint activation of a task phase and location. (f) Level 1 state activations during the last 1000 steps in each environment. Dots are plotted on the agent’s tile and colored by the low-level state. Locations are augmented with random noise for visual clarity. See the main text for further explanation.
@@
Simulated populations of “goal-progress cells” are further illustrated in Figure 4, where their activations are plotted on top of the maze layout during Block 1. The agent’s trajectories are shown as colored dots (with small amounts of noise added to avoid overlapping points). In the third panel, two distinct trajectories can be observed between C and D. The dots are color-coded as before, indicating that along each trajectory (e.g., from A to B in the first panel), different populations of “goal-progress cells” become active in sequence: first those tuned to mid goal expectation (green), then to high goal expectation (yellow), and finally to low goal expectation after the reward is collected at the goal location (purple). Importantly, this ramping pattern emerges consistently across different trajectories, independent of the specific goal destination, while the firing fields of these cells flexibly stretch or contract depending on the physical distance between goals.
A second prominent class of cells reported by (El-Gaby et al., 2024) encode the pursuit of a specific abstract goal (A, B, C, or D), independently of its physical position in the maze (see Figure 2d in (El-Gaby et al., 2024)). In our model, neural activations with this property emerge naturally when considering the agent’s Level 2 (prior) belief over the current task phase (i.e., moving toward A, B, C, or D), which reflects the agent’s expectation about where the next reward will be obtained. Figure 4 illustrates this by plotting the agent’s trajectories during the four Blocks, with colors indicating the current belief over the task phase. The activation corresponding to each phase (e.g., blue when moving from A to B, orange when moving from B to C, green when moving from C to D, and purple when moving from D to A) remains constant along the trajectory until the goal is reached, regardless of the path taken. Crucially, this coding pattern generalizes across Blocks even when goals occupy different spatial locations, thus demonstrating goal-selective but spatially invariant representations—mirroring the empirical findings.
A third prominent class of cells identified by (El-Gaby et al., 2024) encode joint information about the specific abstract goal (A, B, C, or D) and its spatial location in the maze. In our model, such goal-and-location–tuned activations emerge naturally within the grounding likelihood. This is illustrated in Figure 4, where each set of colored dots corresponds to an entry of the grounding likelihood that maps a given spatial location to a task state. This mechanism produces unique activation patterns: activations are identical (or highly similar) only when both the spatial location and the task state coincide, and they differ otherwise. For instance, the activation pattern for location A in Tasks 1 and 4 is identical, whereas the pattern for location A in Tasks 1 and 2 is distinct—mirroring the empirical findings.
@@ 3 Discussion
This study introduces a novel computational approach—schema-based hierarchical active inference (S-HAI)—that addresses the formation of schemas from experience and their rapid generalization to novel contexts. The model builds on theories of hierarchical predictive processing and active inference (Parr et al., 2022; Van de Maele et al., 2024; Pezzulo et al., 2018; Butz et al., 2025; Pezzulo et al., 2018; Smith et al., 2022; Lanillos et al., 2021; Matsumoto and Tani, 2020; Friston et al., 2021; Taniguchi et al., 2022; Isomura and Friston, 2018) and extends them with schema-based mechanisms. S-HAI is organized hierarchically: the higher level (Level 2) is responsible for schema learning and navigation in an abstract task space, while the lower level (Level 1) encodes spatial information and supports navigation in physical space. Crucially, the two levels are connected by a mechanism unique to our model, the grounding likelihood, which maps abstract goal codes in the schema to physical locations. The rapid learning of this mapping enables the agent to flexibly generalize the same schema to novel goal configurations.
@@
With a series of simulations, we validate the ability of the schema-based S-HAI agent to reproduce behavioral findings in tasks requiring rapid generalization, as well as neural findings reported in the medial frontal cortex of rodents performing such tasks. Our results show that, after learning a schema for a general class of navigation problems sharing the same structure—namely, problems requiring cyclic visits to four goal locations to obtain rewards, as in the ABCD task (El-Gaby et al., 2024)—the S-HAI agent exhibits rapid online generalization to novel problems in which the relational structure is preserved but the spatial locations of the four goals change. Remarkably, the schema-based agent outperforms an agent trained offline on all problems, showcasing the advantages of schema-based learning in novel contexts. Our simulations also show that the same approach can be generalized to more challenging tasks in which multiple goals can share the same location (ABCB task), analogous to spatial alternation tasks (Jadhav et al., 2012), and that the S-HAI agent can learn online and select among a mixture of likelihood mappings between the abstract schema and concrete problems, demonstrating the ability to decide when to reuse an existing mapping or create a novel one—capturing at least basic signatures of the assimilation (of novel experiences into existing relational structures) and accommodation (of novel relational structures) processes as conceived by (Piaget, 1952). Finally, and importantly, the S-HAI model reproduces key neural signatures of schemas identified in the medial frontal cortex of rodents performing the ABCD task (El-Gaby et al., 2024), most prominently capturing the activity of goal-progress cells, while also reflecting the heterogeneous coding of other cells sensitive to combinations of goal, spatial, and task-related information.
@@
By providing a mechanistic model of schema-based learning and inference, our framework also generates novel empirical predictions that can be tested in future experiments. One such prediction concerns the behavior and neural representations that might be observed in animals performing the ABCB task (Figure 2). Our simulations suggest that correctly solving this task requires a mechanism (clone-based or similar) capable of disambiguating between distinct instances of the same goal (e.g., goal B). This, in turn, should produce specific patterns of behavior and Level 2 neural representations, including separate representations of the same goal when it is encountered twice (see Figure 1). Another critical prediction concerns the functional role of the neural activity patterns reported in the medial frontal cortex of rodents during the ABCD task (El-Gaby et al., 2024). Our model maps distinct populations of neurons—tuned to goal progress, goal identity, and goal identity in conjunction with spatial location—onto distinct computational processes, namely: goal expectations, beliefs about the current task phase, and the grounding likelihood. As a result, perturbing these neurons should have dissociable effects. For example, disrupting neurons that encode beliefs about the current task phase should impair the animal’s ability to correctly infer its next goal, whereas disrupting neurons that encode the grounding likelihood should impair the animal’s ability to flexibly link task phases to spatial representations. These predictions remain to be tested in future experiments.
@@
The current S-HAI agent has several limitations that can be addressed in future research. First, while our model considers multiple grounding likelihoods, it currently implements only a single schema. The S-HAI framework allows a straightforward extension from maintaining a mixture of grounding likelihoods to also maintaining a mixture over multiple schemas, thereby providing a more comprehensive account of the assimilation and accommodation processes envisaged by (Piaget, 1952). Second, the model focuses primarily on schema learning and inference in the frontal cortex, based on probabilistic generative models that extend Hidden Markov Models (HMMs). This approach permits reproducing some key aspects of the neural coding of schemas in the frontal cortex, but cannot capture its full complexity. Future studies could explore biologically realistic implementations of HMMs (Kappel et al., 2014) and more systematically investigate the mapping between these models and neural computations in the frontal cortex. Additionally, future work might extend S-HAI to provide a system-level model addressing schema-based processes beyond the frontal cortex, encompassing other relevant brain regions such as the hippocampus and entorhinal cortex.

### v2

@@ Abstract
Schemas – abstract relational structures that capture the commonalities across experiences – are thought to underlie humans’ and animals’ ability to rapidly generalize knowledge, rebind new experiences to existing structures, and flexibly adapt behavior across contexts. Despite their central role in cognition, the computational principles and neural mechanisms supporting schema formation and use remain elusive. Here, we introduce schema-based hierarchical active inference (S-HAI), a novel computational framework that combines predictive processing and active inference with schema-based mechanisms. In S-HAI, a higher-level generative model encodes abstract task structure, while a lower-level model encodes spatial navigation, with the two levels linked by a grounding likelihood that maps abstract goals to physical locations. Through a series of simulations, we show that S-HAI reproduces key behavioral signatures of rapid schema-based generalization in spatial navigation tasks, including the ability to flexibly remap abstract schemas onto novel contexts, resolve goal ambiguity, and balance reuse versus accommodation of novel mappings. Crucially, S-HAI also reproduces prominent neural codes reported in rodent medial prefrontal cortex during a schema-dependent navigation and decision task, including task-invariant goal-progress cells, goal-and-spatially conjunctive cells, and place-like codes at the lower level. Taken together, these results provide a mechanistic account of schema-based learning and inference that bridges behavior, neural data, and theory. More broadly, our findings suggest that schema formation and generalization may arise from predictive processing principles implemented hierarchically across cortical and hippocampal circuits, enabling the generalization of experience.
@@ 1 Introduction
Finally, various studies indicate that the frontal cortex—possibly in interaction with the hippocampus—plays a crucial role in forming cognitive maps and schematic associations, integrating new information into existing schemas that support generalization (Gilboa and Marlatte, 2017; Zeithamova et al., 2012; Van Kesteren et al., 2012; Giuliano et al., 2021; Bonasia et al., 2018; Baldassano et al., 2018; Basu et al., 2021; Manakov et al., 2025; Schuck et al., 2016; Wang and Hayden, 2021; Vaidya and Badre, 2022; Samborska et al., 2022). For example, a study in which rodents navigated and generalized rules across distinct environments showed that this ability depends on coordinated hippocampal–prefrontal activity: the hippocampus formed environment-specific representations, whereas prefrontal population activity generalized across environments (Tang et al., 2023).
A particularly striking example of schema learning and generalization is provided by El-Gaby et al. (2024), who demonstrated that the rodent medial frontal cortex is involved in learning the abstract structure of a sequential task—the ABCD task, which requires reaching four goal locations in the correct order—and in reusing this structure in novel environments where the sequence remains the same but the goal locations vary. A key indicator of fast, schema-based learning is that rodents rapidly move toward location A after discovering location D. At the neuronal level, this ability is supported by various types of cells that are sensitive to different combinations of abstract goal, physical location, and other task-relevant information.
@@
While prior work on CSCGs has addressed the problem of learning space as a latent structure from partially observable sensations (Raju et al., 2024), and CSCG schemas (Guntupalli et al., 2023; Swaminathan et al., 2023) addresses the problem of transferring a learned spatial or algorithm structure using schemas, they do not address the problem of transferring multiple task structures that are learned in the same spatial environment. To do this in a partially observable setting, new task-structure schemas have to be learned that use the latent states of the spatial structure that was previously learned.
@@ 2.2 Schema-based hierarchical active inference (S-HAI)
Crucially, the S-HAI agent also includes a grounding likelihood: a probabilistic mapping between the abstract schema representing transitions between goals in task space (i.e., A, B, C, and D) and the concrete locations of the goals in navigation space (i.e., the locations in the maze where it can find reward). Learning grounding likelihoods is what affords schema-based generalization: it enables the agent to rapidly map its abstract schema (e.g., ABCD) to each novel spatial configuration of the goals (i.e., each block), rather than relearning the correct sequence of actions and goals from scratch on each trial. In our simulations, we implement online learning of the grounding likelihood, alongside both an online and offline variant for the Level 2 schema. In addition, we introduce a mixture model over grounding likelihoods that allows the S-HAI agent to flexibly infer which of its existing grounding likelihoods is most useful in the current maze, or to create a new one when needed. Previewing our results, we found that in all cases (online or offline, with or without mixture), grounding likelihoods allow the S-HAI agent to outperform alternative models trained on the same or even larger datasets but without schema learning. Learning the grounding likelihoods is similar to learning the emission matrix of a schema, as in (Guntupalli et al., 2023) and (Swaminathan et al., 2023).
@@
In the hierarchical active inference (HAI) agent without schemas, Level 2 is implemented using a clone-structured graph (CSCG) (George et al., 2021), and the grounding likelihood is an identity matrix. The agent is denoted as HAI-$i$, with the index $i$ indicating the number of tasks the agent is trained on; training is done offline on a sequence of the first $i$ tasks, with 10 000 interaction steps per task. See Figure 1 for the transition dynamics between goals learned by the HAI-$2$ agent (trained on two blocks) at Level 2. Note that, unlike the S-HAI agent, which learns a unique sequential schema, the HAI-$2$ agent learns a unique subcycle for each block. When the HAI-$i$ agent is trained on more blocks, it tends to learn multiple block-specific subcycles (not shown here, but similar to what is reported in (Van de Maele et al., 2024)).
@@
Figure 2 shows the simulation results for the ABCD task, reporting the average reward rate (smoothed over 250 steps), normalized against optimal performance, across 40 blocks. The solid line denotes the mean across blocks, and shaded areas indicate the standard error. Each block ends when the agent reached 10 000 interactions with the environment. As expected, the HAI agents without schemas (trained offline on half (20) or all (40) blocks; HAI-20, and HAI-40) outperformed the Random baseline, with performance improving as the number of training blocks increased. However, the HAI-20 agent trained in half of the blocks showed limited generalization to novel blocks beyond its training set.
@@
Figure 4: Simulation of neural activity in the S-HAI model during the ABCD task. (a) Four blocks of the ABCD task. (b) Simulated activity of “goal-progress” cells during the ABCD task, plotted in a circle. The nodes represent locations visited during the task and their colors nodes indicate the normalized expected inductive cost, under the variational posterior over the Level 1 state; these range from purple (low expectation, corresponding to early phase of progress towards a goal) to green (mid expectation and mid phase of progress) and yellow (high expectation and late phase of progress). The line indicates where the reward was received, and the color indicates which phase of the task this was in. (c) Simulated activity of “goal-progress” cells during the ABCD task, plotted in space. The panels show activations of the normalized expected inductive cost under the variational posterior over Level 1 state in distinct phases of Task 4. As before, purple, green, and yellow represent low expectation (early phase), mid expectation (mid phase), and high expectation (late phase of progress). Note that in the third panel, there are two paths from C to D, i.e., the animal can move upward or to the left after leaving C. (d) Simulated spatial firing fields of the grounding likelihood. Shaded nodes illustrate lagged fields for anchors placed at goal locations in Block 1 (Top) and the resulting remapping in Block 2 (Bottom). This architecture simulates neurons that fire at fixed task-space lags from an anchor. (e) Polar plot depicting simulated neural tuning of three neurons coding for the conjunction of spatial location and progress toward a goal. The right plot indicates the spatial tuning for each of the three neurons, shown in blue (Neuron 1), orange (Neuron 2), and green (Neuron 3). Neurons 1 and 3 are tuned to “mid” progress, and Neuron 2 to “late” progress. (f) Level 1 state activations during the last 1000 steps in each environment. Dots are plotted on the agent’s tile and colored by the low-level state. Locations are augmented with random noise for visual clarity. See the main text for further explanation.
@@
Simulated populations of “goal-progress cells” are further illustrated in Figure 4, where their activations are plotted on top of the maze layout during Block 1. The agent’s trajectories are shown as colored dots (with small amounts of noise added to avoid overlapping points). In the third panel, two distinct trajectories can be observed between C and D. The dots are color-coded as before, indicating that along each trajectory (e.g., from A to B in the first panel), different populations of “goal-progress cells” become active in sequence: first those tuned to mid goal expectation (green), then to high goal expectation (yellow), and finally to low goal expectation after the reward is collected at the goal location (purple). Importantly, this ramping pattern emerges consistently across different trajectories, independent of the specific goal destination. In this simple implementation, goal-progress cells reduce to goal-distance cells, as progress is measured relative to the latent states at Level 1, which map directly onto physical locations. To ensure that the visualized values are comparable across blocks, we normalize the expected inductive cost values.
Another key finding of the ABCD study (El-Gaby et al., 2024) is that neurons do not generalize their task-state preference (A, B, C, or D) across blocks; instead, they remap. However, this remapping is modular: within each module (or “buffer”), one neuron serves as an “anchor” tuned to a specific conjunction of goal progress and location, while other neurons fire at fixed lags in task space relative to that anchor. In this way, the neurons belonging to the same module conserved their within-module tuning relations across tasks.
In our model, the grounding likelihood provides a natural scaffold for the modular remapping of goal codes in the context of planning. Under the active inference framework, actions are selected by scoring the imagined future using the expected free energy. This rollout can be considered “egocentric”—centered on the agent’s current location (e.g., the location of goal A immediately after it has been collected) within the ABCD plan—such that the planning tree encodes multiple expected positions of the agent in task space after n steps (goal B at the next step, goal C after two steps, and so on), all in parallel (Donnarumma et al., 2025). Figure 4 illustrates the simulated neural firing field of the selected high level goal locations, remapped through the grounding likelihood tensor. In each plot, the anchor corresponds to a specific goal location (marked with an X), and the remaining cells encode future goal locations as task-space offsets from this anchor, with increasingly pale shading indicating greater offset.
The first column of Figure 4 illustrates this when location 1 (marked X) is anchored: in Block 1, this location corresponds to goal A, and the module encodes A (anchor), B (anchor+1), C (anchor+2), and D (anchor+3), shown in progressively lighter shades of red. In Block 2, the same location corresponds to goal C, and the module correspondingly remaps to encode C (anchor), D (anchor+1), A (anchor+2), and B (anchor+3). The second column show analogous modular remapping when the anchor is located in position 2. These examples illustrate that while individual neurons remap and change their goal preference across blocks, the internal task-space structure of the module is preserved, consistent with empirical findings and the ”buffers” model of (El-Gaby et al., 2024).
The study of (El-Gaby et al., 2024) further reports that modular remapping is not restricted to goal locations but extends to anchors at intermediate locations between goals, in conjunction with specific, fine-grained goal-progress phases. In our model, this phenomenon can be simulated by anchoring the activity of “goal-progress cells” (shown in Figure 4) to specific spatial locations. Figure 4 illustrates this effect. Here, the activity of goal-progress neurons (discretized into “early,” “mid,” and “late” progress bins) is projected onto a task-centric polar plot. Cardinal angles (0°, 90°, 180°, 270°) represent the four goals of the ABCD loop, while colored neurons are anchored to the spatial locations shown in the adjacent plot. Neuron 1 (blue) and Neuron 2 (orange) are anchored to the same spatial location but encode different progress phases (mid vs. late), whereas Neuron 3 (green) shares mid-progress tuning with Neuron 1 but is anchored to a different location. In this example, neurons appear to remap their task-state preference as goal locations change across Blocks 1–4. However, the anchors—specific conjunctions of progress phase and spatial location—remain invariant across tasks, consistent with the empirical findings of (El-Gaby et al., 2024). For example, the apparent remapping of Neuron 3—from the $A\rightarrow B$ transition in Block 1 to the $C\rightarrow D$ transition in Block 4—reflects the fact that its anchor location corresponds to mid goal-progress between $A$ and $B$ in Block 1, but to mid goal-progress between $C$ and $D$ in Block 4. This result shows that internal task variables—specifically, goal-progress representations—provide a principled computational basis for the modular remapping observed in the study of (El-Gaby et al., 2024).
@@ 3 Discussion
This study introduces a novel computational approach—schema-based hierarchical active inference (S-HAI)—that addresses the formation of schemas from experience and their rapid generalization to novel contexts. The model builds on theories of hierarchical predictive processing and active inference (Parr et al., 2022; Van de Maele et al., 2024; Pezzulo et al., 2018; Butz et al., 2025; Smith et al., 2022; Lanillos et al., 2021; Matsumoto and Tani, 2020; Friston et al., 2021; Taniguchi et al., 2022; Isomura and Friston, 2018) and extends them with schema-based mechanisms. The S-HAI agent is organized hierarchically: the higher level (Level 2) is responsible for schema learning and navigation in an abstract task space, while the lower level (Level 1) encodes spatial information and supports navigation in physical space. Crucially, the two levels are connected by a mechanism unique to our model, the grounding likelihood, which maps abstract goal codes in the schema to physical locations. The rapid learning of this mapping enables the agent to flexibly generalize the same schema to novel goal configurations.
@@
With a series of simulations, we validate the ability of the schema-based S-HAI agent to reproduce behavioral findings in tasks requiring rapid generalization, as well as neural findings reported in the medial frontal cortex of rodents performing such tasks. Our results show that, after learning a schema for a general class of navigation problems sharing the same structure—namely, problems requiring cyclic visits to four goal locations to obtain rewards, as in the ABCD task (El-Gaby et al., 2024)—the S-HAI agent exhibits rapid online generalization to novel problems in which the relational structure is preserved but the spatial locations of the four goals change. remarkably, the schema-based agent outperforms an agent trained offline on all problems, showcasing the advantages of schema-based learning in novel contexts. Our simulations also show that the same approach can be generalized to more challenging tasks in which multiple goals can share the same location (ABCD task), analogous to spatial alternation tasks (Jadhav et al., 2012), and that the S-HAI agent can learn online and select among a mixture of likelihood mappings between the abstract schema and concrete problems, demonstrating the ability to decide when to reuse an existing mapping or create a novel one—capturing at least basic signatures of the assimilation (of novel experiences into existing relational structures) and accommodation (of novel relational structures) processes as conceived by (Piaget, 1952). Finally, and importantly, the S-HAI model reproduces key neural signatures of schemas identified in the medial frontal cortex of rodents performing the ABCD task (El-Gaby et al., 2024), most prominently capturing the activity of goal-progress cells, while also reflecting the heterogeneous coding of other cells sensitive to combinations of goal, spatial, and task-related information.
@@
By providing a mechanistic model of schema-based learning and inference, our framework also generates novel empirical predictions that can be tested in future experiments. One such prediction concerns the behavior and neural representations that might be observed in animals performing the ABCD task (Figure 2). Our simulations suggest that correctly solving this task requires a mechanism (clone-based or similar) capable of disambiguating between distinct instances of the same goal (e.g., goal B). This, in turn, should produce specific patterns of behavior and Level 2 neural representations, including separate representations of the same goal when it is encountered twice (see Figure 1). Another critical prediction concerns the functional role of the neural activity patterns reported in the medial frontal cortex of rodents during the ABCD task (El-Gaby et al., 2024). Our simulations further suggest that distinct components of the model—Level 2 beliefs about goals, Level 1 beliefs about locations, and the grounding likelihood—map onto distinct neuronal populations tuned to goal progress, spatial location, and their conjunctions. Consequently, perturbing these populations should produce dissociable behavioral effects. Disrupting neurons that encode beliefs about the current task phase should impair the animal’s ability to correctly infer its next goal. Disrupting neurons that encode the grounding likelihood should impair systematic remapping across blocks. Finally, disrupting neurons that encode beliefs about spatial location should impair path planning and navigation. These predictions remain to be tested in future experiments.
@@
The current S-HAI agent has several limitations that can be addressed in future research. First, while our model considers multiple grounding likelihoods, it currently implements only a single schema. The S-HAI framework allows a straightforward extension from maintaining a mixture of grounding likelihoods to also maintaining a mixture over multiple schemas, thereby providing a more comprehensive account of the assimilation and accommodation processes envisaged by (Piaget, 1952). Second, the model focuses primarily on schema learning and inference in the frontal cortex, based on probabilistic generative models that extend Hidden Markov Models (HMMs). This approach permits reproducing some key aspects of the neural coding of schemas in the frontal cortex, but cannot capture its full complexity. For example, although Level 2 representations in our model encode goals explicitly, neurons that generalize goal coding across blocks were not observed in the mPFC of the rodent study (El-Gaby et al., 2024). Notably, however, goal-coding neurons that generalize across tasks have been reported in primates (Genovesio et al., 2012). Whether a comparable form of abstract, task-general goal coding exists in the rodent brain remains an open empirical question.
Third, future studies could explore biologically realistic implementations of HMMs (Kappel et al., 2014) and more systematically investigate the mapping between these models, the “structured memory buffer” model of (El-Gaby et al., 2024), and neural computations in the frontal cortex. Additionally, future work might extend S-HAI to provide a system-level model addressing schema-based processes beyond the frontal cortex, encompassing other relevant brain regions such as the hippocampus and entorhinal cortex.